# Bird’s-Eye View to Street-View: A Survey

### v1

@@ II-D Semantic Feature Learning and Representation-based Models
II-D1 PGSL* [5]
: Zhai et al. presented a novel training technique for extracting meaningful features from aerial photos. The aim is to utilize the associated aerial image to predict the ground-level semantic layout. Four networks handle the features, as shown in Figure 5. The CNN network “A” extracts features from the aerial image via VGG16, and then it is fed to PixelNet [38] to convert it to pixel-level. The CNN network, “S”, employs the extracted features from aerial images to control the transformation. The network “F” estimates the change between viewpoints. Ground-level semantic labeling is created by applying the transformation, “T”, to the aerial semantic features. The novelty of this study lies in a new CNN design that links the appearance of an aerial image to a ground-level semantic layout located in the same location. Further, the PGSL [5] method can be extended to ground-level image orientation estimation, synthesis, and localization. Three different initializations have been experimented with: the proposed method initialization, Xavier [39] random initialization, and the VGG16 pre-trained on ImageNet initialization. The results demonstrated a superior performance of the proposed method compared to other initialization techniques on the CVUSA benchmark dataset.
Figure 5: PGSL* [5].
II-D2 WAGL* [2]
: Workman et al. address the wide-area geo-localization (WAGL) task employing pre-existing CNNs to extract the features from ground-level images and then learn to predict those features from the aerial photos with the exact location learning the semantic feature representation jointly. WAGL consists of two training models: a single-scale model and a multi-scale model. The first model primarily aims to extract aerial features corresponding to ground-level images. The second model matches ground-level images to aerial photos at multiple spatial scales. The novelty lies in the in-depth analysis of existing CNN architectures for cross-view localization tasks with cross-view training for learning semantic features jointly from multiple image sources. The results showed that the features learned through multi-scale training outperformed single-scale and other methods [40, 41] on Charleston and San Francisco benchmark datasets.
II-E Geometric Correspondence Models
@@
II-E1 SSVS* [4]
: It is a multi-task image synthesizing and retrieval architecture. SSVS stands for satellite-to-street view synthesis for geo-localization. The network learns the latent representations that are helpful for retrieval if they are employed for generating satellite and street-view images. In the proposed network, the generator is designed as U-Net [10], which consists of residual blocks, while the discriminator is constructed as a PatchGAN classifier [23]. The retrieval network consists of two main components, the encoder block and the spatial attention module, which convert an image’s local features into global representation, as illustrated in Figure 6(a). The results showed that the proposed method outperformed other models such as Pix2Pix [23], X-Fork [11], X-Seq [11], and PGSL [5] on CVUSA and CVACT datasets.
@@ II-E2 SSVG* [7]
: SSVG (Satellite and Street View Geometry) generates street-view images from satellite views based on modeling the geometric correspondences. Shi et al. [7] explicitly states the geometric correspondences between the satellite and street-view images, easing the information transfer between the mentioned domains. The proposed model applied the Satellite to Street-View Image Projection (S2SP) module to determine a geometric correspondence between street-view and satellite images. First, given the satellite image, the height probability is estimated through Pix2Pix [23], which is used for height estimation. Next, a satellite-view multiplane image (MPI) is constructed, given the satellite-view image and the estimated height probability distribution. Stretching and unrolling occur on each concrete cylinder to obtain a consistent street view. From street-view MPI, the final street-view image is generated in back-to-front order. Results are presented on two benchmark datasets: CVUSA [2] and CVACT [18] and illustrated that the SSVG [7] outperformed Pix2Pix [23] and X-Fork [11] on both the datasets.
II-E3 JLO [15]
@@
: A recently proposed joint localization and orientation estimation method by Shi et al. [15] is based on Dynamic Similarity Matching (DSM) to measure the feature similarity of the image pair. The authors handle the cross-view domain gap by using polar transform to align the aerial image up to an unknown azimuth angle, which helps the network learn the feature correspondences between polar aerial and ground images instead of having to learn the geometry relation between these two domains. Figure 6(b) shows an overview of the JLO method, where the polar transform is applied to an aerial image. Two stream-CNN networks are applied to extract polar-transformed aerial and ground-level image features. The orientation of the ground-level image concerning the aerial image is estimated using the correlation given the extracted feature volume representation. After that, the aerial features are shifted and cropped to get the image part that probably matches ground features, and the similarity is used for location retrieval. The novelty of this method is that it presents the first image-based geo-localization technique for estimating the orientation and position of ground-level images independent of their field of view. The JLO achieved the highest top-1 accuracy compared with other recent works [2, 5, 19, 17, 18, 14], with a value of 99.67% for CVUSA [2] and 97.32% for CVACT [18].
II-E4 LONN* [18]
@@
stands for Lending Orientation to Neural Networks is a ground-level to aerial image matching method based on a Siamese network with explicit orientation encoding of each image pixel. The proposed method has been evaluated using the CVUSA benchmark dataset. In this study, Siamese-type two-branch CNN has been applied, where each branch will learn the useful features for comparing the views. The understanding of an image’s orientation facilitates the localization task, and based on that; the authors presented a method for injecting the orientation information for each pixel via orientation maps. Figure 6(c) shows an overview of the LONN [18], where ground-level and aerial images are fed to the Siamese network along with their associated orientation maps. The triplet loss function receives the two learned feature vectors and uses them to train the network. The orientation map injection can be done in two ways: either concatenated with the input layer only or injected in an intermediary convolution layer. The novelty of this work lies in proposing a simple approach for incorporating the orientation information of each pixel. Also, a novel Siamese-based CNN network is used to learn the feature embedding orientation and appearance information. The results showed that the proposed method outperformed other recent works Workman et al.  [2], Zhai et al.  [5], Vo et al. [19], and Hu et al.  [17] with recall top-1 equal to 93.19%.
II-E5 GT-CrossView [19]
@@
: This model tackled ground-level to satellite image matching by investigating two categories of CNN networks. The first one is representation learning networks for embedding the cross-view images into the same feature space and classification networks for identifying matches. The second category is Siamese and triplet networks for computing similarity. A new Distance-Based Logistic Loss (DBL) function has introduced a novel version of Siamese and triplet networks. In this study, the orientation alignment is considered in addition to the spatial one, so learning the rotation invariant (RI) representation of satellite images helps learn complex representations without using more or larger filters and is done by simply applying random rotation to the training data, i.e., data augmentation. For better learning of representation, the authors proposed an auxiliary loss function where the added rotation amount can be utilized as a label. Figure 6(d) illustrates the network architecture that augments data through random rotation and incorporates an additional branch for orientation regression. The results demonstrate that both classification and hybrid (Siamese classification) networks performed better than feature representation networks, with an accuracy of 90% and 91.5%, respectively. Also, Siamese-DBL and triplet-DBL outperformed other original networks with accuracy rates of 88% and 90.2%, respectively. The images of a street view panorama are obtained by randomly querying the Google Map of the United States across 11 cities.
II-E6 CVM-Net [17]
@@
: It is an architecture for cross-view ground-to-aerial geolocation, consisting of two subnetworks, as shown in Figure 6(e). The first one, CVM-Net-I, contains two aligned NetVALDs layers that collect the local features into one common space from various views. The second one, CVM-Net-II, consists of two weight-shared NetVALDs layers to convert the local features into a shared space before aggregating to get the global descriptor. CVM-Net [17] is notable because the authors combine a Siamese network with NetVALDs layers and learn robust cross-view image representations for matching tasks. In addition, the authors proposed a novel loss called weighted soft margin, which enhances retrieval accuracy and makes the training convergence faster. The results illustrate that CVM-Net-I outperforms CVM-Net-II when AlexNet and VGG-16 are employed as local feature extractors; the accuracy reaches 65.4% and 91.4%, respectively. Furthermore, CVM-Net-I and CVM-Net-II outperform other existing methods [5][19][2], and the Siamese network of VGG16 and AlexNet baselines, achieving an accuracy of 67.9% for cropped images and 91.4% for panorama images evaluated on CVUSA and Dayton [19], respectively.
II-F Siamese Network-based Models
@@
II-F1 Where-CNN [16]
: An early Siamese network [42] inspired the street-view and aerial image matching technique known as Where-CNN. It consists of two CNNs modified from [43] (see Figure 7) having different or identical standard parameters. Regarding different parameters, both learn domain-specific deep representations (called Where-CNN-DS). On the other hand, when parameters are shared, both learn a general representation between aerial and street-view images (called Where-CNN). Moreover, the contrastive foss function proposed in [44] is utilized and is trained using Caffe [45]. The features extracted via Where-CNN and Where-CNN-DS outperform Image-CNN, Places-CNN, and hand-crafted features. Further, Where-CNN-DS, a domain-specific model, achieved slightly better performance than Where-CNN, a domain-independent model. The street view and 45-degree aerial images are collected from Google across seven cities.
Figure 7: Lin et al. architecture overview [16].
II-G MLP-based Models
@@ II-G1 CrossMLP [9]
: It is a two-stage framework for aerial-to-street-view generation with a novel cascaded Cross MLP-Mixer (CrossMLP) sub-network as the first stage and a refined pixel loss as the second stage. Figure 8 depicts the overall architecture, wherein the first stage, the Cross-MLP sub-network, stacks multiple Cross-MLP blocks for generating the street view and its corresponding semantic map progressively. Each CrossMLP module captures the structure and appearance information within each view. It can learn the latent correspondence between two inputs by applying a sequential MLP. Then, latent cues guide the transformation from aerial to street view. The refined pixel loss helps with the noisy semantic label by adding more reasonable regularization to the total optimization loss, and with less redundancy, the pixel-level loss becomes more compact. The novelty of this work lies in proposing a two-stage framework for generating street-view images using a novel CrossMLP module. Using CrossMLP to learn the latent mapping cues between aerial-view and street-view semantic maps helps with cross-view translation with accurate appearance and geometry structure. The results showed that CrossMLP performed better than other recent works: Pix2Pix [23], X-Fork [11], X-Seq [11], and SelectionGAN [13], except in terms of PSNR, SD, and SSIM, achieved a bit lower results compared with the others. This study evaluates the method on two benchmark datasets: CVUSA [2] and Dayton [19].
Figure 8: Ren et al. architecture overview [9].
TABLE I: Quantitative comparisons on the CVUSA dataset.
Accuracy (%) Inception Score
Method Top-1 Top-5 All Top-1 Top-5 SSIM PSNR SD KL
Pix2Pix [23] 7.33 9.25 25.81 32.67 3.2771 2.2219 3.4312 0.3923 17.6578 18.523 59.81 ± 2.12
X-Seq [11] 20.58 31.24 50.51 63.66 3.4432 3.5567 3.5567 0.4356 19.0509 18.67 11.71 ± 1.55
X-Fork [11] 15.98 24.14 42.91 54.41 3.8151 2.6738 4.0077 0.4231 18.8067 18.438 15.52 ± 1.73
SelectionGAN [13] 41.52 65.51 74.32 89.66 3.8074 3.9197 3.9197 0.5323 23.1466 19.61 2.96 ± 0.97
CrossMLP [9] 44.96 69.96 76.98 91.91 3.8392 2.8035 3.9757 0.5251 23.1532 19.58 2.69 ± 0.94
Real Data 7 - - - 4.8741 3.2959 4.9943 - - - -

### v2

@@ 2.4 Representation-based Models
PGSL* [57]: Zhai et al. presented a novel training technique for extracting meaningful features from aerial photos. The aim is to utilize the associated aerial image to predict the ground-level semantic layout. Four networks handle the features, as shown in Figure 6. The CNN network “A” extracts features from the aerial image via VGG16, and then it is fed to PixelNet [2] to convert it to pixel-level. The CNN network, “S”, employs the extracted features from aerial images to control the transformation. The network “F” estimates the change between viewpoints. Ground-level semantic labeling is created by applying the transformation, “T”, to the aerial semantic features. The novelty of this study lies in a new CNN design that links the appearance of an aerial image to a ground-level semantic layout located in the same location. Further, the PGSL [57] method can be extended to ground-level image orientation estimation, synthesis, and localization. Three different initializations have been experimented with: the proposed method initialization, Xavier [14] random initialization, and the VGG16 pre-trained on ImageNet initialization. The results demonstrated a superior performance of the proposed method compared to other initialization techniques on the CVUSA benchmark dataset.
Figure 6: PGSL: Semantic Feature Learning and Representation-based Models [57].
WAGL* [53]: Workman et al. address the wide-area geo-localization (WAGL) task employing pre-existing CNNs to extract the features from ground-level images and then learn to predict those features from the aerial photos with the exact location learning the semantic feature representation jointly. WAGL consists of two training models: a single-scale model and a multi-scale model. The first model primarily aims to extract aerial features corresponding to ground-level images. The second model matches ground-level images to aerial photos at multiple spatial scales. The novelty lies in the in-depth analysis of existing CNN architectures for cross-view localization tasks with cross-view training for learning semantic features jointly from multiple image sources. The results showed that the features learned through multi-scale training outperformed single-scale and other methods [52, 60] on Charleston and San Francisco benchmark datasets.
2.5 Geometric Correspondence Models
@@
SSVS* [48]: It is a multi-task image synthesizing and retrieval architecture. SSVS stands for satellite-to-street view synthesis for geo-localization. The network learns the latent representations that are helpful for retrieval if they are employed for generating satellite and street-view images. In the proposed network, the generator is designed as U-Net [39], which consists of residual blocks, while the discriminator is constructed as a PatchGAN classifier [18]. The retrieval network consists of two main components, the encoder block and the spatial attention module, which convert an image’s local features into global representation, as illustrated in Figure 11(a). The results showed that the proposed method outperformed other models such as Pix2Pix [18], X-Fork [36], X-Seq [36], and PGSL [57] on CVUSA dataset.
@@
SSVG* [42]: SSVG (Satellite and Street View Geometry) generates street-view images from satellite views based on modeling the geometric correspondences. Shi et al. [42] explicitly states the geometric correspondences between the satellite and street-view images, easing the information transfer between the mentioned domains. The proposed model applied the Satellite to Street-View Image Projection (S2SP) module to determine a geometric correspondence between street-view and satellite images. First, given the satellite image, the height probability is estimated through Pix2Pix [18], which is used for height estimation. Next, a satellite-view multiplane image (MPI) is constructed, given the satellite-view image and the estimated height probability distribution. Stretching and unrolling occur on each concrete cylinder to obtain a consistent street view. From street-view MPI, the final street-view image is generated in back-to-front order. Results are presented on two benchmark datasets: CVUSA [53] and CVACT [30] and illustrated that the SSVG [42] outperformed Pix2Pix [18] and X-Fork [36] on both the datasets.
@@
JLO [43]: A recently proposed joint localization and orientation estimation method by Shi et al. [43] is based on Dynamic Similarity Matching (DSM) to measure the feature similarity of the image pair. The authors handle the cross-view domain gap by using polar transform to align the aerial image up to an unknown azimuth angle, which helps the network learn the feature correspondences between polar aerial and ground images instead of having to learn the geometry relation between these two domains. Figure 11(b) shows an overview of the JLO method, where the polar transform is applied to an aerial image. Two stream-CNN networks are applied to extract polar-transformed aerial and ground-level image features. The orientation of the ground-level image concerning the aerial image is estimated using the correlation given the extracted feature volume representation. After that, the aerial features are shifted and cropped to get the image part that probably matches ground features, and the similarity is used for location retrieval. The novelty of this method is that it presents the first image-based geo-localization technique for estimating the orientation and position of ground-level images independent of their field of view. The JLO achieved the highest top-1 accuracy compared with other recent works [53, 57, 50, 16, 30, 37], with a value of 99.67% for CVUSA [53] and 97.32% for CVACT [30].
@@
LONN* [30]: stands for Lending Orientation to Neural Networks is a ground-level to aerial image matching method based on a Siamese network with explicit orientation encoding of each image pixel. The proposed method has been evaluated using the CVUSA benchmark dataset. In this study, Siamese-type two-branch CNN has been applied, where each branch will learn the useful features for comparing the views. The understanding of an image’s orientation facilitates the localization task, and based on that; the authors presented a method for injecting the orientation information for each pixel via orientation maps. Figure 11(c) shows an overview of the LONN [30], where ground-level and aerial images are fed to the Siamese network along with their associated orientation maps. The triplet loss function receives the two learned feature vectors and uses them to train the network. The orientation map injection can be done in two ways: either concatenated with the input layer only or injected in an intermediary convolution layer. The novelty of this work lies in proposing a simple approach for incorporating the orientation information of each pixel. Also, a novel Siamese-based CNN network is used to learn the feature embedding orientation and appearance information. The results showed that the proposed method outperformed other recent works Workman et al. [53], Zhai et al.  [57], Vo et al. [50], and Hu et al.  [16] with recall top-1 equal to 93.19%.
@@
GT-CrossView [50]: This model tackled ground-level to satellite image matching by investigating two categories of CNN networks. The first one is representation learning networks for embedding the cross-view images into the same feature space and classification networks for identifying matches. The second category is Siamese and triplet networks for computing similarity. A new Distance-Based Logistic Loss (DBL) function has introduced a novel version of Siamese and triplet networks. In this study, the orientation alignment is considered in addition to the spatial one, so learning the rotation invariant (RI) representation of satellite images helps learn complex representations without using more or larger filters and is done by simply applying random rotation to the training data, i.e., data augmentation. For better learning of representation, the authors proposed an auxiliary loss function where the added rotation amount can be utilized as a label. Figure 11(d) illustrates the network architecture that augments data through random rotation and incorporates an additional branch for orientation regression. The results demonstrate that both classification and hybrid (Siamese classification) networks performed better than feature representation networks, with an accuracy of 90% and 91.5%, respectively. Also, Siamese-DBL and triplet-DBL outperformed other original networks with accuracy rates of 88% and 90.2%, respectively. The images of a street view panorama are obtained by randomly querying the Google Map of the United States across 11 cities.
@@
CVM-Net [16]: It is an architecture for cross-view ground-to-aerial geolocation, consisting of two subnetworks, as shown in Figure 11(e). The first one, CVM-Net-I, contains two aligned NetVALDs layers that collect the local features into one common space from various views. The second one, CVM-Net-II, consists of two weight-shared NetVALDs layers to convert the local features into a shared space before aggregating to get the global descriptor. CVM-Net [16] is notable because the authors combine a Siamese network with NetVALDs layers and learn robust cross-view image representations for matching tasks. In addition, the authors proposed a novel loss called weighted soft margin, which enhances retrieval accuracy and makes the training convergence faster. The results illustrate that CVM-Net-I outperforms CVM-Net-II when AlexNet and VGG-16 are employed as local feature extractors; the accuracy reaches 65.4% and 91.4%, respectively. Furthermore, CVM-Net-I and CVM-Net-II outperform other existing methods [57][50][53], and the Siamese network of VGG16 and AlexNet baselines, achieving an accuracy of 67.9% for cropped images and 91.4% for panorama images evaluated on CVUSA and Dayton [50], respectively.
2.6 Siamese Network-based Models
@@
Where-CNN [29]: An early Siamese network [8] inspired the street-view and aerial image matching technique known as Where-CNN. It consists of two CNNs modified from [23] (see Figure 8) having different or identical standard parameters. Regarding different parameters, both learn domain-specific deep representations (called Where-CNN-DS). On the other hand, when parameters are shared, both learn a general representation between aerial and street-view images (called Where-CNN). Moreover, the contrastive foss function proposed in [15] is utilized and is trained using Caffe [19]. The features extracted via Where-CNN and Where-CNN-DS outperform Image-CNN, Places-CNN, and hand-crafted features. Further, Where-CNN-DS, a domain-specific model, achieved slightly better performance than Where-CNN, a domain-independent model. The street view and 45-degree aerial images are collected from Google across seven cities.
Figure 8: WhereCNN [29] architecture overview.
2.7 MLP-based Models
@@
CrossMLP [38]: It is a two-stage framework for aerial-to-street-view generation with a novel cascaded Cross MLP-Mixer (CrossMLP) sub-network as the first stage and a refined pixel loss as the second stage. Figure 9 depicts the overall architecture, wherein the first stage, the Cross-MLP sub-network, stacks multiple Cross-MLP blocks for generating the street view and its corresponding semantic map progressively. Each CrossMLP module captures the structure and appearance information within each view. It can learn the latent correspondence between two inputs by applying a sequential MLP. Then, latent cues guide the transformation from aerial to street view. The refined pixel loss helps with the noisy semantic label by adding more reasonable regularization to the total optimization loss, and with less redundancy, the pixel-level loss becomes more compact. The novelty of this work lies in proposing a two-stage framework for generating street-view images using a novel CrossMLP module. Using CrossMLP to learn the latent mapping cues between aerial-view and street-view semantic maps helps with cross-view translation with accurate appearance and geometry structure. The results showed that CrossMLP performed better than other recent works: Pix2Pix [18], X-Fork [36], X-Seq [36], and SelectionGAN [46], except in terms of PSNR, SD, and SSIM, achieved a bit lower results compared with the others. This study evaluates the method on two benchmark datasets: CVUSA [53] and Dayton [50].
Figure 9: CrossMLP architecture overview taken from [38].
2.8 Recent Advancements
BEVGen [45]: It proposed a conditional GAN-based model for generating realistic street-view images from Bird’s Eye View layout for traffic scenarios. The evaluation is conducted on the NuScenes 360 driving dataset [7]. Figure 10 (a) shows an overview of BEVGen architecture, where it encoded both street-view images and BEV layout to latent discrete representation and flattened them for the autoregressive transformer. Subsequently, the spatial embeddings are added to the street view and Birds Eye View (BEV) tokens, and the pairwise bias is combined with the attention weight. After that, the decoder receives the tokens to get the generated street-view images. BEVGen compares with the baseline model, which has the same GPT architecture and uses identical first-stage encoders and decoders. The result indicates that BEVGen outperformed the baseline on metrics such as Fréchet Inception Distance (FID), road mIoU, and vehicle mIoU. As a limitation, BEVGen [45] faces challenges in accurately generating small objects like people and specific vehicles.
Sat2Scene [28]: Li et al. introduces framework for direct 3D scenes video generation given satellite images based on diffusion model. 3D sparse representations have been incorporated with the diffusion model to ensure the images’ consistency and enable the generation of street-view features that closely associated with the geometry directly in 3D space. Sat2Scene consists of three steps: generation step, feature extraction step, and rendering step. In the generation step assigns color for the foreground point cloud by utilizing 3D diffusion model combined with the sparse convolutions, while background panorama is generated using a 2D diffusion model. The feature extraction step is responsible for extracting scene features closely associated with the point cloud. The rendering step generates scene frames from arbitrary views using neural rendering. Figure 10 (b) shows an overview of Sat2Scene architecture. The proposed model was evaluated on two datasets: OmniCity [27] and HoliCity [61] datasets. The results demonstrated that, Sat2Scene is able to generate a realistic and texture-consistent street-view image sequences.
CrossViewDiff [26]: Lin et al. proposed a diffusion-based framework for satellite to street-view synthesis task. The CrossViewDiff model presented a novel cross-view texture and satellite scene structure estimation modules, as illustrated in Figure 10 (c). It leveraging from the imaging and geometric relationships between the satellite and street-view pairs to extract the textual and structural information from the satellite images. These extracted information have been used as a control to guide the street-view generation process, resulting in improved fidelity of the generated street-view images. The authors have used GPT-4o for evaluating the generated street-view images, in addition to the standard evaluation metrics, providing more comprehensive evaluation of the generated street-view images. The proposed model was evaluated on three datasets: CVUSA, CVACT, and OmniCity [27]. The results lustrated that, CrossViewDiff achieves better results with an average increase of $39\%$ in FID, $35.5\%$ GPT-based score, and $9\%$ in SSIM, outperforming other models.
Refer to caption Refer to caption
(a) (b)
Figure 10: Recent Advancements . (a) BevGen [45], (b) Sat2Scene [28], (c) CrossViewDiff [26]
After reviewing the existing methods for satellite to street-view synthesis task, we can see that GANs have been widely adopted in this domain. However, GANs may suffer from several issues during the training process, which results in instable training and affect the generated results. The first one is mode collapse which results to a lack of diversity in the generated results. The second one is vanishing gradient issue which arises when the discriminator is very effective in classifying the images into real and fake, which results in training instability and no feedback provided to the generator for improvements. The the third one is non-convergence issue which occurs due to gradient optimizer’s failure to reach Nash equilibrium. These issues in during GAN training, resulting in generated images with unrealistic, blurry, and less diversity outputs [40]. Nonetheless, the recent advancements, transformers and diffusion models offer significant improvements that help to overcome these limitations. Diffusion models it works by gradually removing Gaussian noise from images to generate an image that closely reflect the target distribution, which improves the training stability and results in the generation of high-quality, more realistic, and finer details. However, diffusion models are computationally expensive, it requires several iterative steps in denoising process, resulting in increasing the time and memory usage which limit their scalability for real-time applications and large datasets. On the other hand, transformers shows strong capabilities in modeling long-range spatial dependencies, enabling them to capture detailed global context, which is considered as important feature for synthesizing detailed environment, such as street-view images [25]. However, it is computationally expensive and memory intensive which affect its scalability in limited resources environments.
Refer to caption Refer to caption
(a) (b)
Refer to caption Refer to caption
(c) (d)
Study Year Dataset Method Task Metrics Pros and Cons
WAGL [53] 2015 - Charleston - San Francisco - CNN - Cross-view matching - Accuracy Pros:
- Applying satellite features extraction at multiple spatial scales enhances the method’s ability to handle perspectives variations, resulting in improved the performance of cross-view matching.
Cons:
- The multi-scale feature extraction from satellite images introduces computational complexity which may limit its salability for real-world applications with limited resources.
Where-CNN [29] 2015 - Collected from Google across seven cities. - CNN - Siamese network - Cross-view matching - Accuracy Pros:
- The method performs geo-localization using only satellite image, eliminating the need for street-view images, which make it applicable in scenarios where the ground-level data unavailable.
Cons:
- The orientation estimation of the street-view images could be unknown, the system need to test different possible orientation during the testing time to find the correct orientation. This process could be time-consuming and computationally expensive, making it less practical for time-sensitive applications.
GT-CrossView [50] 2016 - CVACT - CNN - Siamese network - Triplet network - Cross-view matching
- Accuracy
- Recall
Pros:
- The explicit prediction of relative rotations of satellite and street-view images improves the retrieval accuracy.
Cons:
- Sensitive to the quality of rotation predication. Incorrect rotation can negatively impact the retrieval accuracy.
PGSL [57] 2017 - CVUSA - CNN - GAN - Cross-view matching - Cross-view synthesis - Average precision Pros:
- The method demonstrates the ability of extract semantic features directly from the satellite images, therefore avoiding the need of manual annotation.
Cons:
- The use of multiple sub-networks increases the computational and memory requirements.
X-Seq, X-Fork [36] 2018 - CVUSA - Dayton - cGAN - Cross-view synthesis
- Accuracy
- Inception score
- SSIM
- PSNR
- Sharpness
Difference
Pros:
- Both X-Fork and X-Seq generate both street-view images and their corresponding semantic segmentation maps, facilitating the learning of more informative features that preserve the street-view consistency.
Cons:
- the reliance on annotated semantic segmentation data, limits the scalability of the propose methods to datasets that lack such annotations.
CVM-Net [16] 2018 - CVUSA - Dayton - CNN - VLAD descriptor - Cross-view matching - Recall Pros:
- Efficient training by utilizing a weighted soft margin loss that accelerate the training convergence.
- CVM-Net-I not sharing weights between the two views NetVLAD branches, helps the network to learn view-specific representations, resulting in more effective learning for cross-view matching task.
Cons:
- Non-shared NetVLADs restults increase the number of parameters, resulting in higher model complexity.
cGANs [12] 2018 - Collected from Geograph API and Google Map static API - cGAN - Cross-view synthesis - Land-cover classification - Accuracy Pros:
- The method explores and compares different embedding techniques, demonstrating its flexibility and provides insight into how they affect the generated results and help for future enhancements.
Cons:
- Despite using dense feature maps, the generated images lack details, limiting the method effectiveness for downstream tasks such as classification.
Table 1: Summary of Reviewed Related Studies in the Literature
Study Year Dataset Method Task Metrics Pros and Cons
BridgeGAN [65] 2019 - Collected by [23] - GAN - Cross-view synthesis
- SSIM
- PSNR
- LPIPS
Pros:
- Utilizing homograph view works as bridge between the frontal and bird-eye views, enabling model to learn the transformation gradually.
Cons:
- The use of three distinct GANS for separate domain frontal, homography and bird-eye views, with shared layers, makes the model computationally expensive during training phase.
SelectionGAN [46] 2019 - Dayton - CVUSA - Ego2Top - cGAN - Cross-view synthesis
- Accuracy
- Inception score
- SSIM
- PSNR
- SD
Pros:
- Multi-channel attention module enhances the learning effectiveness.
- Cascaded semantic guidance generation, allow the model to generate detailed images by progressively refining the initial generation.
Cons:
- The multi-channel attention and multi intermediate generation introduces more computational cost.
FeatureFusion [37] 2019 - CVUSA - cGAN - Cross-view matching - Cross-view synthesis - Accuracy Pros:
- Incorporating edge maps and semantic segmentation provides additional boundary information.
- Applying feature fusion strategy helps to integrate features from different source, resulting in enhancing the feature learning.
- Multi-resolution aggregation from multiple convolution layers allow the model to preserve global context to local details representations.
Cons:
- Feature fusion and multi-resolution aggregation introduce computational complexity.
LONN [30] 2019 - CVUSA - CVACT - CNN - Siamese network - Cross-view matching - Recall Pros:
- Incorporating per-pixel orientation information in a straightforward manner through orientation maps, allow to inject the geometric information without increasing model complexity.
- The simple design of the model contributes to facilitating model training, debug and deployment, improving its usability for real-world applications.
Cons:
- The simplicity of this design limits its scalability to more complex geospatial tasks, which require more deep feature modeling.
LGGAN [47] 2020 - Dayton - CVUSA - GAN - Cross-view synthesis
- Inception score
- Accuracy
- SSIM
- PSNR
- KL
- SD
Pros:
- Semantic guided generation, which ensures the structural consistency of the generated images.
- Combines global and local class-specific generation, enabling the model to maintain the overall scene coherence and capture local details.
Cons:
- Multiple interconnected components increased the overall model complexity, making training more challenging.
GeoUrban [32] 2020 - Collected from Google Street-view using Google API - GAN - Cross-view synthesis
- Accuracy
- PSNR
- SSIM
- SD
- mIoU
Pros:
- Employ pre-processing step to mitigate the image misalignment by calculating overlap ratio.
Cons:
- Does not correct the misaligned data, it just filter out images with low ratio which may affect the dataset diversity and model robustness.
JLO [43] 2020 - CVUSA - CVACT - CNN - Dynamic similarity matching - Cross-view matching - Recall Pros:
- Robust to image misorientation and narrow Field of View (FoV).
- Effective orientation estimation using azimuth angle.
Cons:
- Azimuth angle sensitivity, where small misalignment affect the localization accuracy.
Table 1: Summary of Reviewed Related Studies in the Literature (continued)
Study Year Dataset Method Task Metrics Pros and Cons
SSVS [48] 2021 - CVUSA - CVACT - cGAN - ResNet34 - Cross-view matching - Cross-view synthesis
- Recall
- SSIM
- PSNR
- SD
- KL-Scores
- LPIPS
Pros: Multi-tasking framework allow shared features learning.
Cons: Join learning increases the computational complexity.
CrossMLP [38] 2021 - Dayton - CVUSA - GAN - Cross-view synthesis
- Accuracy
- Inception score
- SSIM
- PSNR
- SD
- KL
Pros:
- Addresses the noise in the input semantic segmentation, improving the model robustness against noise.
Cons:
- The multi-stage architecture, along with components like feature fusion and Cross-MLP blocks, introduces architectural complexity and resource-intensive model.
SSVG [42] 2022 - CVUSA - CVACT - cGAN - Cross-view synthesis
- Accuracy
- RMSE
- SSIM
- PSNR
- SD
- mIoU
Pros:
- Applies implicit height map supervision leveraging two-view geometric constraints, thus avoiding the need for annotated data, which may not always available.
Cons:
- Sensitive to height estimation errors, leading to large difference between the generated image and its corresponding ground truth when estimating is incorrect.
PAGAN [54] 2022 - CVUSA - Orlando- Pittsburgh (OP) cGAN Cross-view synthesis
- Accuracy
- SSIM
- PSNR
- SD
- KL
- FID
Pros:
- Cross-stage Attention module provide efficient feature fusion by using lightweight attention mechanism.
- Orientation-aware data augmentation, making the model robust for views changes.
Cons:
- The lightweight attention may limit the model capabilities in capturing complex features.
PanoGAN[54] 2022 - CVUSA - Orlando- Pittsburgh (OP) - cGAN - Cross-view synthesis
- Inception Score
- Accuracy
- SSIM
- PSNR
- SD
- KL
Pros:
- Iterative image refinement.
- Effective multi-scale feature representation that capture both high-level and low-level information with minimal computation cost.
Cons:
- Resources-intensive due to multiple components.
- Architecture complexity.
- Dependency on labeled data, segmentation map, which may not always available.
BevGen[45] 2024 - NuScenes - Autoregressive Transformer - Cross-view synthesis - FID scores Pros:
- Multi-view generation.
- Spatial consistency modeling.
Cons:
- Multi-view generation increases model complexity.
CrossViewDiff [26] 2024 - CVUSA - CVACT - OmniCity - Diffusion model - Cross-view synthesis
- SSIM
- PSNR
- SD
- FID
- KID
Pros:
- Incorporating structure-aware and texture mapping modules, improving the layout consistency and accurate appearance for the generated street-view images.
Cons:
- Integrating diffusion model with additional guidance modules introduces high computational costs, as diffusion by is inherently resource-intensive by nature.
Sat2Scene [28] 2024 - OmniCity - HoliCity - Diffusion model - Cross-view synthesis
- SSIM
- PSNR
- LPIPS
- FID
- KID
- FVD
- KVD
Pros:
- 3D street-view generation, is valuable features for real-world applications, like urban planning simulation.
Cons:
- Combining diffusion models with 3D sparse representations, may introduce computational and memory cost, especially with large dataset.
Table 1: Summary of Reviewed Related Studies in the Literature (continued)
3 Datasets